Crew Photos – April 15th

Photo of the Day: Crew practicing towing an ATV after a medical emergency simulation EVA.


Operations Report – April 15th

Crew 192 Operations Report 15Apr2018

SOL 7

Name of person filing report: Richard Blakeman

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 84%

Diesel – 75% fuel

Propane – 64% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Approximately 4 gallons left in one and the other is empty

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 815 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133514.4

Toilet tank emptied: Yes. Emptied today, 15 April.

ATV’s Used: Blue ATVs #1, #2

Oil Added? No but checked and OK.

ATV Fuel Used: 5%

# Hours the ATVs were Used today: 0.5

Notes on ATVs: All working well and ready for use, gas filled up today, all at 100%.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 38.2
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: A quiet day at the hab.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, two one-piece suits and two two-piece suits used. Radios working fine. Spacesuit 9 ventilation wasn’t working yesterday, but worked today. May have just needed a charge/charging port may not have been inserted all the way. All good.

Summary of Hab operations: Stand-down day, crew relaxed around the hab. Also took the opportunity to perform some training. See commander’s report.

Summary of GreenHab operations: Plants watered twice.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank prepped for additional cleaning in the RAM.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

EVA Report – April 15th

EVA Report

EVA #9 – 15Apr2018 – Sol 7

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Andreea Radulescu, Victoria Varone

Objectives:

To simulate a medical emergency requiring evacuation while on an EVA. Additional objective is to tow an ATV using an ATV.

Operations:
1230-1500

Narrative: Two crew members went out on EVA and called in a medical emergency. The doctor was dispatched and an extra crew member stayed IVO the HAB to support the single crew member left inside the HAB or the doctor as required given a 5 person crew.

Upon RTB we faked an ATV with a problem which could not be fixed. The crew the practiced towing the ATV into the compound.

Vehicles
Blue ATVs #1 and #2, Rover Spirit

EVA Report – April 15th

EVA Report

EVA #9 – 15Apr2018 – Sol 7

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Andreea Radulescu, Victoria Varone

Objectives:

To simulate a medical emergency requiring evacuation while on an EVA. Additional objective is to tow an ATV using an ATV.

Operations:
1230-1500

Narrative: Two crew members went out on EVA and called in a medical emergency. The doctor was dispatched and an extra crew member stayed IVO the HAB to support the single crew member left inside the HAB or the doctor as required given a 5 person crew.

Upon RTB we faked an ATV with a problem which could not be fixed. The crew the practiced towing the ATV into the compound.

Vehicles
Blue ATVs #1 and #3, Rover Spirit

Sol Summary – April 15th

Crew 192 Sol Summary Report 15Apr2018

Sol 7

Summary Title: It’s Not Possible, It’s Necessary

Author: Victoria Varone

Mission Status: Today was a half-stand down day for the crew.

Sol Activity Summary: One medical emergency simulation EVA in the afternoon, which taught the crew the basics of how to handle a medical emergency when out in the field on Mars. Crew doctor Ashok Narayanamoorthi shared the rules of what to do and what not to do, the most important being to remain calm.

The afternoon consisted of a debrief, some knot-tying lessons with the Commander, a Martian Matinee (we watched "Interstellar"), and some more fun cooking for dinner.

Look Ahead Plan: We’re aiming for some solid EVAs this week, exploring the stretches of the Martian terrain.

Anomalies in Work: None

Weather: Warm, partly sunny early in the day, cloudy in the afternoon.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report

Support Requested: None

Maintenance request: Almost out of bleach, could use a refill.

Commander Report – April 15th

Terrific day of rest and training combined. Crew slept in and the pancakes were a big hit. We held extensive knot training on how and where to tie knots. The crew was maybe a little too enthused about tying things up… ;*)…..they all progressed to a single handed bowline rapidly. A simple life saving knot.

IMPRESSED: Great training and plan on the part of our Doc. (Blue 02- Ashok) He planned and executed a superb medical training plan without a hitch and everyone learned something.

LEARNED: Lessons from training include and are not limited to…reduce communication during work to that which is imperative because people need two hands to drive or work and it is difficult to respond when working if not essential. We need to do knot training regularly and we will every day. When towing an ATV heavy person should be on operating vehicle. We all learned….SLOW IS SMOOTH …AND SMOOTH IS FAST.

~Dituri sends

Warm Regards,

Joseph Dituri, PhD(c)
CDR, U.S. Navy Diving Officer (ret)

Crew Photos – April 14th